# Feedforward suppression of readout-induced faults in quantum error correction

### v1

@@ Abstract
We propose a method to reduce readout-induced faults, applicable in settings like quantum error correction with repeated measurement cycles. The method consists of an adaptive readout sequence conditioned on each check qubit’s readout result from the previous cycle. For readout errors and measurement-induced leakage that are stronger in a particular qubit state, this feedforward protocol can suppress the physical qubit errors. Focusing on a simple realization of conditionally flipping (by an X gate) the state of check qubits before their measurement, we investigate the effect of such state-dependent errors using simulations in the setup of a low-density parity check code. We show that the suggested protocol can reduce both logical errors and decoding time, two important aspects of fault-tolerant quantum computations.
@@
Ingredients of quantum error correction (QEC) and fault-tolerant computing are being realized with various platforms [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. A dominant approach in the field is based on quantum circuits realizing stabilizer codes, wherein errors in data qubits (which store the quantum information) are repeatedly mapped to check qubits that are then measured to extract the error syndromes [13]. Such codes therefore share a significant reliance on mid-circuit readout of check qubits. A measurement-based scheme [14, 15] also forms the basis of efficient implementations of logical gates on state-of-the-art quantum low-density parity check (LDPC) codes [16, 17, 18, 19, 20, 21]. Measurement results (syndromes) are fed to a decoder whose task is to determine the changes in the state of the data qubits, and specify a possible correction procedure.
@@
However, instead of applying an actual correction, the Pauli frame (including the logical one) can be tracked by the decoder [22, 23]. Pauli frame tracking is often required since decoding is a computationally intensive task, while solid-state qubits operate at relatively fast clock times [24, 25]. For example, superconducting two-qubit gates require an order of a hundred nanoseconds or even less, and syndrome cycles require about one microsecond. This makes decoding syndrome data in real time an additional outstanding challenge for practical QEC implementations. Approaches to reduce decoding time are therefore an active area of research from both algorithmic and hardware perspectives [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38].
Figure 1: A schematic depiction of the adaptive readout sequence, based on outcome of each qubit’s previous measurement. (a) As an example focused on two check qubits of a stabilizer-based error correction code, following the measurement step shown on the left and in the absence of any further error, the first and second check qubits are again expected to be measured in the states |$\ket{1}$ and |$\ket{0}$ respectively (at the end of the syndrome cycle “SC”). The measurement type (indicated by color) applied to each qubit is conditioned and adapted to its expected state. (b) A possible realization of the adapted measurement is a conditional flip (an X gate) applied to each qubit that is expected to be in |$\ket{1}$; in this example, setting the first qubit back to |$\ket{0}$ while the second one is left untouched. A classical NOT on the first qubit’s measured bit would recover the expected syndrome at the end.
Over many syndrome cycles with Pauli frame tracking the fraction of check qubits measured in the excited state is expected to saturate at one half. For example, in CSS-type codes, this is the case for X checks starting a cycle with all qubits in the ground state due to random collapse, and for the Z checks this follows from a random walk on the syndromes. A check qubit’s value at the end of the syndrome cycle will tend to repeat its last value until a fault leads to a change in the check. The large fraction of excited check qubits can potentially lead to increased errors, since the excited state is often more prone to errors than the ground state. Importantly, leakage of the qubit state from the computational space (or qubit loss) is a significant issue, and readout-induced leakage may be stronger from one particular state, in particular the excited state [39, 40, 41, 42, 43, 44, 45]. Various approaches to active and passive leakage reduction form an intense area of investigation [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. Efforts in the domain of readout-error mitigation are centered on improving accuracy given a large number of shots [59, 60, 61, 62, 63, 64], with Ref. [65] improving sampling results with X gates added before the readout of qubits that are likely to be in the higher-error state at the termination. Alternatively, works addressing single-shot improvements, which are relevant for common QEC contexts, require additional qubits that complicates the device connectivity [66, 67].
@@
In this paper we propose an adaptive gate approach to reduce faults in the error-prone state during mid-circuit readout of a series of identical circuits, e.g., the syndrome cycles in a QEC setup. The protocol is designed to suppress the underlying physical cause of readout, preparation and leakage errors by using simple-to-implement dynamic circuits with a bit-wise feedfoward (conditional) operation, which naturally fits within a Pauli frame-tracking scenario. In addition to the reduction in logical error rates, decoding becomes faster as fewer detectors are unsatisfied.
@@
The adaptive readout method. The essence of the method, as schematically depicted in Fig. 1(a), is in adapting the measurement gate of each check qubit to its expected state at the end of every QEC syndrome cycle. The adapted measurement gate sequence is selected based on the readout result from the previous cycle. More generally, the conditional logic could be extended to depend on a majority vote or some other processing of the last $m$ readout results. The state-adapted measurement gate sequence can include, e.g., tuning the readout pulse parameters (amplitude, frequency, and other hardware specific parameters), or the discrimination boundaries and algorithm differentiating between the states 0 and 1, allowing an overall optimization of the error rate. Such adaptations can be advantageous in particular for competing errors (e.g., when leakage is stronger from one qubit state while readout error is biased towards the other state). In a similar scenario, these adaptations could potentially be used to reduce some measurement-induced errors from one state at the cost of increasing others (e.g., decreasing leakage at the cost of increasing readout errors), even if these errors are not biased originally, as discussed later.
@@
In the following we focus on the adaptation by insertion of a conditional gate before the measurement pulse. Without loss of generality, we choose that the excited state $|1\rangle$ suffers from higher rates of readout-induced faults than the ground state $|0\rangle$. Then a conditional local X gate can be applied to each check qubit based on its last readout bit, as in Fig. 1(b). Check qubits that have been correctly measured in the state $|1\rangle$ in the previous syndrome cycle, are now expected to be measured in the ground state $|0\rangle$ when applying the adaptive gate before the measurement (i.e., without the adaptive gate and assuming Pauli frame tracking, the check would otherwise be again mapped to $|1\rangle$). A classical bit flip is applied to the adaptive gate outcome, keeping the original meaning for the syndrome as if there was no adaptive gate. The reduced frequency of measured $|1\rangle$ is then expected to reduce the overall number of errors. Qubits whose measurement result was 1 due to a readout error will be flipped erroneously, but this will not propagate beyond one more cycle. In a syndrome cycle implementation that avoids resets of the check qubits, the protocol should be conditioned on the previous two measurements – applied if 00 was measured in the previous cycle and $1$ was measured in the one before it In the rest of this work we provide numerical evidence indicating that this protocol could reduce logical error rates and decoding times, in particular accounting for leakage errors.
@@
Decoding model. We simulate the LDPC high-encoding-rate Gross code with 144 data qubits and the same number of check qubits (encoding 12 logical qubits), which is described in detail in [68]. The Gross code is a CSS code for which we simulate noisy syndrome measurement circuits for both X and Z checks. A syndrome cycle (a Clifford circuit), which can be thought of as a logical idle gate, is composed of single-qubit gates of different types and two-qubit CNOT gates coupling data and check qubits. The single-qubit gates include preparation (qubit initialization) and readout (measurement) gates of the check qubits, and some idle (identity) gates on the data qubits. We execute Monte Carlo simulations of repeated applications of the syndrome cycle together with random realizations of Pauli errors occurring following every gate within it with prescribed probabilities. The decoding simulations are based on a standard circuit-level noise model as implemented in [69]. We must stay within the boundaries of Pauli noise (or very close to a Pauli model [7]) as it remains the only feasible option at those scales despite its known limitations [70, 71, 72].
We focus on error parameters consistent with current state-of-the-art superconducting qubits, and our goal in the numerical simulations is to explore the overall trends in the parameter space, which necessarily come with some tradeoff between various choices. The standard noise model includes errors in the preparation and the readout (wherein an error is a state flip), in the idle gates (a depolarizing single-qubit Pauli channel), and in the CNOT gate (a depolarizing two-qubit Pauli channel).
@@
We extend this model with state-dependent errors of readout, readout-induced leakage, and further relevant error types as detailed below. To account for readout-induced leakage, we consider a model within which only check qubits in the state $|1\rangle$ can leak outside of the computational levels, and this happens during readout. A leaked qubit always reads 1 and has a seepage probability to return to the state $|1\rangle$ during ensuing readouts. We assume that the state of a leaked check does not change during the CNOT gates of the syndrome cycle. Within this model, a CNOT gate between a data qubit and a leaked check qubit has some probability of incurring (independently) a Pauli Z or X error to the coupled data qubit, which we denote as the backaction error rate. This is a plausible reduction of the standard two-qubit Pauli error of a CNOT gate.
@@
To simulate state-dependent errors with different rates as in Eq. (1), we start every simulation with a random initial state of the data qubits, in order to induce a random Pauli frame. Since the competing leakage and seepage (with different rates) lead to a steady state in the population of leaked qubits, we study in Appendix A its characteristics and we start each simulation with a random initialization of leaked qubits corresponding to this mean population.
In addition, we consider two types of idle (identity) gates on data qubits; short, simultaneous with a CNOT gate, and long (with an increased associated error), simultaneous with the readout gates on the check qubits. We use uniform qubit error rates, fixing the preparation, short idle and CNOT error parameters;
@@
The gate parameters reflect values expected to be achieved in early systems running large-scale QEC, while the readout error rates are on the higher side of mean values in large superconducting quantum processors. As exemplified by our results in the following, it may be advantageous to accommodate higher readout errors in order to lower other significant readout-induced faults (especially leakage), as discussed below. In the data below we will vary the leakage, seepage, backaction and long idle error rates. In the simulation of the qubit flip protocol, we test the effect of an increased long-idle error (representing a tradeoff accounting for the longer duration due to the X gate and possibly some delays due to the classical conditioning), and we also account for a flip error in the X gate.
@@
Leakage rates on the order of $10^{-3}$ have been reported with transmon qubits [3] and we vary the leakage rate in this regime. Regarding the seepage rate, in the simulations it is immediately clear (and has been known experimentally) that if leaked qubits are not reset to the computational space fast enough, a leaked population quickly builds up to make error correction impractical. Therefore some form of leakage removal must be assumed, and we vary seepage rates in the regime of order 1. A high seepage rate (corresponding to a frequent application of leakage removal throughout the device) must clearly incorporate some cost to it (in operation time at least), a tradeoff that we explore below.
@@
The logical error rate is also very sensitive to errors on the data qubits being induced by leaked check qubits during CNOT gates. We see that if this backaction probability is taken naively to be of order 1 (corresponding to gates between leaked qubits being blindly executed), the logical error rates rise to a detrimental level. Therefore, this scenario is uninteresting to simulate, and we focus on the regime in which this probability is a few percent, which can be thought of as modeling a mitigation scheme such as identifying the leaked qubits (with a high probability) in real time [73, 74, 75, 76] and avoiding gates involving those qubits. It is expected and we confirm that the higher the leakage and backaction error rates are, and the lower the seepage rate is, the more beneficial the flip protocol will be.
@@
Results. The data presented in the rest of this paper is based on simulating multiple “shots” (independent simulation runs) for each combination of parameters, decoding windows of $N=12$ syndrome cycles starting with the belief propagation (BP) algorithm [26]. BP is an iterative approach to syndrome decoding, and the mean executed number of iterations is observed to grow with the complexity of the error. A cutoff of a maximal number of iterations is imposed at 200 iterations. If BP does not converge, a secondary decoder (with a significantly higher algorithmic complexity) is invoked, which is the ordered statistics decoding (OSD) algorithm of order 7 with combination sweep as [77, 68]. The decoder is based on a linearization of the noise model, searching for the most likely linear combination of the syndromes of single faults, with state-independent error probabilities defined by Eqs. (2)-(3) together with the long-idle error rate $p_{\rm id,l}=0.005$. For $N$ syndrome cycles in a shot of the simulation, if $p_{N}$ is the probability of a logical error in the decoding, the logical error per syndrome cycle is customarily [68] defined as $1-(1-p_{N})^{1/N}\approx p_{N}/N$, and is reported below together with other decoding characteristics.
@@
Figure 2: Results from decoding of simulated syndrome measurements as a function of the backaction error probability (corresponding to leaked check qubits inducing independent Z and X errors on the coupled data qubits), with the data comparing the flip protocol vs. its absence, with error rates as in Eqs. (2)-(5). (a) The mean number of BP iterations in each decoding execution. The decoding time is significantly reduced with the flip protocol due to fewer BP iterations being needed. (b) The logical error rate per syndrome cycle, showing a significant reduction to a nearly constant level. See the text for a detailed discussion.
@@
and scan the magnitude of the backaction error across a small range. For generating syndromes with the flip protocol we increase the long idle gate error, and add also the flip error parameter;
@@
Nevertheless, even in that scenario, the flip protocol would improve significantly the decoding time, an improvement that becomes even more pronounced for increased error rates. This can be deduced from the reduction in the mean number of BP iterations seen in Fig. 2 and moreover, due to the fraction of converged iterations increasing significantly with the flip protocol [not directly shown in Fig. 2 but a similar behavior can be seen in Fig. 3 below]. This is consistent with the large increase in the mean ground-state population of check qubits just before readout, which rises from 50% to approximately 93% for all data points, and the related quantity of the population of leaked check qubits just before readout [see Appendix A]. The number of BP iterations is plausibly reduced due to less detectors being unsatisfied, and the better convergence of BP leads to a large reduction of decoding time following the fewer (costly) secondary decoders executions.
@@
As can be seen in Fig. 2, the decoding performance is sensitive to backaction from leaked qubits, and significantly improves with the flip protocol. A related parameter scan is presented in Fig. 3 where we take an example of an even higher seepage rate
@@
while the backaction probability is fixed to the maximal value shown in the scan of Fig. 2.
@@
Scanning the leakage rate in a relevant range we see again an expected sensitivity in the logical error rate to the leakage probability, and a strong suppression of the error with the flip protocol. We note again that here as well, even at very low leakage, the flip protocol reduces the expected decoding time, with a lower mean number of BP iterations and a higher fraction of converged ones.
In a typical hardware implementation of readout it can be expected that varying the duration of the readout step would incur some tradeoff between different error rates. To investigate this aspect we show in Fig. 4 the decoding results when varying together the seepage rate and the long-idle error (neglecting here a separate increase of the idle error when the flip protocol is performed). A non-monotonic error in the decoding penalties is observed without a flip protocol, exemplifying how such an approach can serve as a guideline to finding optimal operational parameters. By taking advantage of the flip protocol, for the presented parameters the global lowest logical error and decoding time are obtained together when – perhaps surprisingly – the lowest seepage rate is chosen. Combining numerical input with direct experimental optimization of complex tradeoff choices in hardware parameters, software aspects and decoding model can be expected to become increasingly more relevant, encompassing not only logical errors [78, 79] but other computational aspects such as decoding time, as above.
@@
Figure 4: Decoding simulations varying simultaneously the seepage rate and the error in the long-idle gate [with the other error rates as in Eqs. (2)-(3)]. (a) The mean number of BP iterations. (b) The logical error per syndrome cycle. Modeling a tradeoff between the efficacy of a leakage removal process and its cost in terms of the prolonged idling of data qubits, the non-monotonic behavior of the decoding quantities without the flip emphasizes the importance of having a variety of tools at hand for optimizing complementing approaches.
@@
Conclusion and outlook. The suggested protocol is expected to have the highest impact if the conditional or adapted gates on the check qubits could be implemented without delays, and indeed in principle the whole duration of a syndrome cycle can be used to process the conditional execution and prepare the required instruction.
@@
Analyzing the expected tradeoffs between different mechanisms for suppression of errors in QEC settings could be a powerful tool to test global optimization approaches. The simulation code used to produce the data for this work is available as open source [80], and it can be extended and used to quantify more elaborate error models or conditional protocols. As another example for state-dependent noise rates, in Appendix B we probe preparation errors correlated to readout errors (i.e., a preparation error occurring conditioned on a readout error, as in the so-called “conditional reset” scheme), seeing a dramatic increase in the logical error rates under certain conditions, with some modest reduction using the flip protocol.
As mentioned already, it can happen that for the “standard” calibration of the readout pulse, leakage is stronger from one qubit state while the readout error is larger for the other state. In such a case, the readout gate can be tailored to each state such that the readout errors and induced faults will be reduced (with some tradeoff between different types of such faults). Each adaptation will have some effect and the overall error rates can be optimized. This approach might be beneficial even if the “standard” errors were symmetric. To give a concrete example that is simple to analyze, we can focus on readout errors alone and consider a favorable modification of the discrimination. We assume a standard calibration of the readout resulted for each qubit in a minimal mean readout error [$p_{\rm read}$ of Eq. (1)] and some given value of $p_{\rm bias}$ that depends on the qubit and measurement properties. A simple adaptation of the readout would be to set a different threshold of discrimination, decreasing $p(1|0)$ while paying with $p(0|1)$. Defining the prior probability $p_{0}$ of a check qubit to be in the 0 state just before readout at the end of a syndrome cycle, the results in Appendix A show that with the flip protocol we get $p_{0}\approx 0.93$ (for the parameters described there). Then it is clear that by minimizing $p_{0}\times p(1|0)+p_{1}\times p(0|1)$ in the setting of the QEC syndrome cycle, the overall readout error rate can be reduced. Since $p_{0}$ depends on all errors during the syndrome cycle, the best optimization would require an experiment applying the protocol to known states and adapting the measurement for each qubit such that a minimal overall error (or decoding time) is observed.

### v3

@@ Abstract
Qubit measurements in quantum devices involve various types of errors, including erroneous state determination, correlated preparation errors and measurement-induced leakage from the computational states. We propose a feedforward protocol to reduce readout-induced faults, applicable for qubits with errors biased between the different states, in settings like quantum error correction with repeated measurement cycles. The method consists of an adaptive readout sequence conditioned on each check qubit’s readout result from the previous cycle, which is optimized for the expected measured state. Focusing on a simple realization of conditionally flipping (by an X gate) the state of check qubits before their measurement, we investigate the effect of such state-dependent errors using simulations in the setup of a low-density parity check code. We show that the suggested protocol can reduce both logical errors and decoding time, two important aspects of fault-tolerant quantum computations.
@@
Components of quantum error correction (QEC) and fault-tolerant computing are being realized with various platforms [andersen2020repeated, satzinger2021realizing, sundaresan2023demonstrating, PhysRevX.11.041058, krinner2022realizing, postler2022demonstration, google2023suppressing, gupta2024encoding, bluvstein2023logical, google2025quantum, putterman2025hardware, reichardt2024logical]. A dominant approach in the field is based on quantum circuits realizing stabilizer codes, wherein errors in data qubits (which store the quantum information) are repeatedly mapped to check qubits that are then measured to extract the error syndromes [RevModPhys.87.307]. A measurement-based scheme [gottesman2014fault, breuckmann2021quantum] also forms the basis of efficient implementations of logical gates on state-of-the-art quantum low-density parity check (LDPC) codes [cohen2022low, cross2024improved, williamson2024low, swaroop2024universal, cowtan2025parallel, zhang2025time]. Such codes therefore share a significant reliance on the performance of mid-circuit measurements of check qubits.
@@
Measurement results of check qubits constitute syndromes that are fed to a decoder whose task is to determine the changes in the state of the data qubits, and specify a possible correction procedure. Instead of applying an actual correction, the Pauli frame (including the logical one) can be tracked by the decoder [knill2005quantum, Chamberland2018faulttolerant]. Pauli frame tracking is often required since decoding is a computationally intensive task, while solid-state qubits, for example, operate at relatively fast clock times [kjaergaard2020superconducting, RevModPhys.95.025003]. Approaches to reduce the real-time decoding computational burden are therefore an active area of research from both algorithmic and hardware perspectives [roffe_decoding_2020, liyanage2023scalable, bombin2023modular, caune2024demonstrating, gong2024toward, chen2025improved, jones2024improved, kaufmann2025blockbp, demarti2024decoding, higgott2025sparse, PhysRevLett.133.240602, ott2025decision, beni2025tesseract].
[…]
In typical implementations, a check qubit’s value will tend to repeat itself at the end of each syndrome cycle until a fault leads to a flip. The fraction of check qubits measured in the excited state is expected to saturate at one half over many syndrome cycles when using Pauli frame tracking in a system with a computational basis formed by an effective ground state and excited state. For example, in Calderbank, Shor, Steane (CSS)-type codes, this is the case for X checks starting a cycle with all qubits in the ground state due to random collapse, and for the Z checks this follows from a random walk on the syndromes.
With many qubit modalities, error rates are asymmetric between the computational $|0\rangle$ and $|1\rangle$ states during measurement and other gate operations. Importantly, leakage of the qubit state from the computational space (or qubit loss) is a significant issue, and readout-induced leakage may be stronger from one particular state, often the excited state [PhysRevLett.117.190503, PhysRevApplied.20.054008, nesterov2024measurement, dumas2024measurement, PhysRevA.111.012611, PhysRevLett.134.100601, bista2025readout]. Choices in the implementation of QEC could be customized to minimize exposure to the higher error rate states, significantly improving performance, and complementing various hardware approaches to enhance active and passive leakage reduction [PRXQuantum.2.030314, mcewen2021removing, miao2023overcoming, lacroix2025fast, kim2025fast, thorbeck2024high, PhysRevLett.130.250602, camps2024leakage, yang2024coupler, fowler2013coping, ghosh2015leakage, brown2019leakage, suchara2015leakage].
In the domain of readout-error mitigation efforts are centered on increasing accuracy given a large number of shots [PRXQuantum.2.040326, PhysRevA.106.012423, doi:10.1126/sciadv.abi8009, PRXQuantum.5.030315, PRXQuantum.6.010307, henao2023adaptive, koh2024readout, trex], with Ref. [PhysRevA.103.022407] improving sampling results with X gates added before the readout of qubits that are likely to be in the state with a higher error at the termination of the circuit. Those approaches are not necessarily practical for QEC and single-shot improvements are required. Current works addressing such improvements use additional qubits, which complicates the required device connectivity [PhysRevA.105.012419, PhysRevLett.134.080801]. Therefore, new methods for error suppression are needed.
@@
In this paper we propose an adaptive gate approach to reduce faults in the error-prone state during mid-circuit readouts that are repeated as part of a series of identical circuits, e.g., the syndrome cycles in a QEC setup. The protocol is designed to suppress the underlying physical cause of readout, preparation and leakage errors by using dynamic circuits with a feedfoward (conditional) operation, which naturally fits within a Pauli frame-tracking scenario. In addition to the reduction in logical error rates, we find that practical decoding algorithms execute faster as fewer detectors are unsatisfied.
@@
The adaptive readout method. The essence of the method, as schematically depicted in Fig. 1(a), is in adapting the measurement gate of each check qubit to its expected state at the end of every QEC syndrome cycle. The adapted measurement gate sequence is selected based on the readout result from the previous cycle. More generally, the conditional logic could be extended to depend on a majority vote or some other processing of the last $m$ readout results. The state-adapted measurement gate sequence can include, e.g., tuning the readout pulse parameters (amplitude, frequency, and other hardware specific parameters), or the discrimination boundaries and algorithm differentiating between the qubit states, allowing an overall optimization of the error rate. Such adaptations can be advantageous in particular for competing errors (e.g., when leakage is stronger from one qubit state while readout error is higher in the other state). Similarly, these adaptations could potentially be used to reduce some measurement-induced errors from one state at the cost of increasing others (e.g., decreasing leakage while increasing readout errors), even if these errors are not biased originally, as discussed in App. A.
@@
In the following we focus on the adaptation by insertion of a conditional gate before the measurement pulse. Without loss of generality, we choose that the excited state $|1\rangle$ suffers from higher rates of readout-induced faults than the ground state $|0\rangle$. Readout is restricted to this basis and measurement of other axes is facilitated by single qubit operations (e.g., Hadamard gates) before the measurement. Then a conditional local X gate can be applied to each check qubit based on its last readout bit, as in Fig. 1(b). Check qubits that have been correctly measured in the state $|1\rangle$ in the previous syndrome cycle, are now expected to be measured in the ground state $|0\rangle$ when applying the adaptive gate before the measurement (i.e., without the adaptive gate and assuming Pauli frame tracking, the check would otherwise be again mapped to $|1\rangle$). A classical bit flip is applied to the measurement outcome, keeping the original meaning for the syndrome as if there was no adaptive gate. The reduced frequency of measured $|1\rangle$ is then expected to reduce the overall number of errors. Qubits whose measurement result was 1 due to a readout error will be flipped erroneously, but this will not propagate beyond one more cycle. In a syndrome cycle implementation that avoids resets of the check qubits, the protocol should be conditioned on the previous two measurements – applied if 0 was measured in the previous cycle and $1$ was measured in the one before it.
This method can be applied to various error correction codes in order to suppress readout-induced faults. In the rest of this work we focus in detail on a specific code, the LDPC Gross code [bravyi2024high], where we provide numerical evidence indicating that this protocol could reduce logical error rates and decoding times, in particular accounting for leakage errors.
@@
Decoding model. We simulate the LDPC Gross code with 144 data qubits and the same number of check qubits (encoding 12 logical qubits), which is described in detail in [bravyi2024high]. The Gross code is a CSS code and X-type and Z-type syndromes are decoded separately. The decoding simulations are based on a standard circuit-level noise model as implemented in [bravyidecoder]. We execute Monte Carlo simulations of repeated applications of the syndrome cycle together with random realizations of Pauli errors with prescribed probabilities following every gate. We focus on error parameters consistent with near-term state-of-the-art superconducting qubits, and our goal in the numerical simulations is to explore the overall trends in the parameter space, which necessarily come with some tradeoff between various choices. We must stay within the boundaries of stochastic noise (very close to a Pauli model [google2023suppressing]) as it remains the only feasible option at those scales despite its known limitations [greenbaum2017modeling, iverson2020coherence, PhysRevA.111.022613, malekakhlagh2025efficient].
The syndrome cycle is a Clifford circuit, which can be thought of as a logical identity circuit, composed of single-qubit gates of different types and two-qubit CNOT gates coupling data and check qubits. The single-qubit gates include preparation (qubit initialization) and readout (measurement) gates of the check qubits, and some idle (identity) gates on the data qubits. The standard noise model includes errors in the preparation and the readout (wherein an error is a state flip), in the idle gates (a depolarizing single-qubit Pauli channel), and in the CNOT gate (a depolarizing two-qubit Pauli channel).
@@
We extend this model with state-dependent errors of readout, readout-induced leakage, and further relevant error types as detailed below. To account for readout-induced leakage, we consider a simplified model in which only check qubits in the state $|1\rangle$ can leak outside of the computational levels, and leakage is calculated only during readout (as readout-induced faults is the focus of this work). A leaked qubit always reads 1 and its state does not change during the CNOT gates of the syndrome cycle. Within this model, a CNOT gate between a data qubit and a leaked check qubit has some probability of incurring (independently) a Pauli Z or X error to the data qubit, which we denote as the backaction error rate. This is a plausible reduction of the standard two-qubit Pauli error. A leaked qubit has also a seepage rate, which is its probability to return to the state $|1\rangle$ during one syndrome cycle. In our simplified model the seepage is calculated only at the end of each cycle, during readout gate. This is an effective averaged description of spontaneous and induced processes that take the qubit back to the computational space with a prescribed mean rate.
@@
To simulate state-dependent errors with different rates as in Eq. (1), we start every simulation with a random initial state of the data qubits, in order to induce a random Pauli frame. Moreover, since the competing leakage and seepage (with different rates) lead to a steady state in the population of leaked qubits, we first study its characteristics (see App. C), and in the following we start each simulation with a random initialization of leaked qubits corresponding to this mean population.
In addition, we consider two types of idle (identity) gates on data qubits; short, simultaneous with a CNOT gate (denoted $p_{\rm id,\,s}$), and long ($p_{\rm id,\,l}$, with an increased associated error), simultaneous with the readout gates on the check qubits. We use uniform qubit error rates, fixing the preparation, short idle and CNOT error parameters (see Tab. 1 for a summary of all parameters);
@@
The gate parameters reflect values expected to be achieved in early systems running large-scale QEC, while the readout error rates are on the higher side of mean values in large superconducting quantum processors, as the mid-circuit measurements are expected to be faster and more error prone than the reported terminal readouts. As exemplified by our results in the following, it may be advantageous to accommodate higher readout errors in order to lower other significant readout-induced faults (especially leakage or the data qubits’ idle time). In the data below we vary the leakage, seepage, backaction and long idle error rates. In the simulation of the qubit flip protocol, we test the effect of an increased long-idle error (representing a tradeoff accounting for the longer duration due to the X gate and possibly some delays due to the classical conditioning), and we also account for a flip error in the X gate.
@@
Leakage rates on the order of $10^{-3}$ have been reported with transmon qubits [sundaresan2023demonstrating] and we vary the leakage rate in this regime. Regarding the seepage rate, in the simulations it is immediately clear (and has been known experimentally) that if leaked qubits are not reset to the computational space fast enough, a leaked population quickly builds up to make error correction impractical. Therefore some form of leakage removal must be assumed, and we vary seepage rates in the regime of order 1. A high seepage rate (corresponding to a frequent application of leakage removal throughout the device) must clearly incorporate some cost to it (in operation time at least), a tradeoff that we explore below.
@@
The logical error rate is also very sensitive to errors on the data qubits being induced by leaked check qubits during CNOT gates. This type of error is highly dependent on the details of qubit and gate operations, however, we see that if this backaction probability is taken naively to be of order 1 (corresponding to gates involving leaked qubits being blindly executed), the logical error rates rise to a detrimental level. Therefore, this scenario is uninteresting to simulate, and we focus on the regime in which this probability is a few percent, which may reflect qubit design or can be thought of as modeling mitigation schemes that assume identifying the leaked qubits (with a high probability) in real time [bultink2020protecting, varbanov2020leakage, kanazawa2023qutrit, mude2025efficient]. It is expected and confirmed that the higher the leakage and backaction error rates are, and the lower the seepage rate is, the more beneficial the flip protocol will be.
@@
Results. The data presented in the rest of this paper is based on simulating multiple “shots” (independent simulation runs) for each combination of parameters, decoding windows of $N=12$ syndrome cycles starting with the belief propagation (BP) algorithm [roffe_decoding_2020]. BP is an iterative approach to syndrome decoding, and the mean executed number of iterations is observed to grow with the complexity of the error. A cutoff of a maximal number of iterations is imposed at 200 iterations. If BP does not converge, a secondary decoder (with a significantly higher algorithmic complexity) is invoked, which is the ordered statistics decoding (OSD) algorithm of order 7 with combination sweep as in [Roffe_LDPC_Python_tools_2022, bravyi2024high]. The decoder is based on a linearization of the noise model, searching for the (estimated) most likely linear combination of the syndromes of single faults, with state-independent error probabilities defined by Eqs. (2)-(3) together with the long-idle error rate $p_{\rm id,l}=0.005$. For $N$ syndrome cycles in a run of the simulation, if $p_{N}$ is the probability of a logical error on one or more of the encoded qubits, the logical error per syndrome cycle is customarily [bravyi2024high] defined as $1-(1-p_{N})^{1/N}\approx p_{N}/N$, and is reported below together with other decoding characteristics.
@@
Figure 2: Results from decoding of simulated syndrome measurements as a function of the backaction error probability (corresponding to leaked check qubits inducing independent Z and X errors on the coupled data qubits), with the data comparing the flip protocol (denoted as “w/ flip”) vs. its absence (denoted as “w/o flip”), for error rates as in Eqs. (2)-(5). (a) The mean number of BP iterations in each decoding execution. The decoding time is significantly reduced with the flip protocol due to fewer BP iterations being needed. (b) The logical error rate per syndrome cycle, showing a significant reduction to a nearly constant level. See the text for a detailed discussion.
@@
and scan the magnitude of the backaction error across a small range (in the regime of a few percent, as explained previously). For generating syndromes with the flip protocol we increase the long idle gate error, and add also the flip error parameter;
@@
Nevertheless, even in that scenario, the flip protocol would improve significantly the decoding time, an improvement that becomes even more pronounced for increased error rates. This is seen in the reduction in the mean number of BP iterations seen in Fig. 2 and furthermore, because of the fraction of converged iterations increasing significantly with the flip protocol [not directly shown in Fig. 2 but a similar behavior can be seen in Fig. 3 described below]. This is consistent with the large increase in the mean ground-state population of check qubits just before readout, which rises from 50% to approximately 93% for all data points, and the related population of leaked check qubits just before readout (see App. C). The number of BP iterations is plausibly reduced due to less detectors being unsatisfied, and the better convergence of BP leads to a large reduction of decoding time following the fewer (costly) secondary decoders executions.
@@
A related parameter scan is presented in Fig. 3 where we take an example of an even higher seepage rate
@@
while the backaction probability is fixed to the maximal value shown in the scan of Fig. 2. Scanning the leakage rate in a relevant range we see again an expected sensitivity in the logical error rate to the leakage probability, and a strong suppression of the error with the flip protocol. We note again that here as well, even at very low leakage, the flip protocol reduces the expected decoding time, with a lower mean number of BP iterations and a higher fraction of converged ones.
@@
Figure 4: Decoding results when varying the seepage rate (between 0.2 and 0.7 in increments of magnitude 0.1) simultaneously with the error in the long-idle gate (between 0.005 and 0.01 in steps of 0.001), with the other error rates as in Eqs. (2)-(3). (a) The mean number of BP iterations. (b) The logical error per syndrome cycle. Modeling a tradeoff between the efficacy of a leakage removal process and its cost in terms of the prolonged idling of data qubits, the non-monotonic behavior of the decoding quantities without the flip emphasizes the importance of having a variety of tools at hand for optimizing complementing approaches.
In a typical hardware implementation of readout it can be expected that varying the duration of the readout step would incur some tradeoff between different error rates. Specifically, increasing the seepage rate could be expected to require increasing the long idle duration, resulting in an increased error probability on the idling data qubits. To investigate this aspect we show in Fig. 4 the decoding results when varying together the seepage rate and the long-idle error (neglecting here a separate increase of the idle error when the flip protocol is performed). A non-monotonic behavior of the decoding penalties is observed without a flip protocol, exemplifying how such an approach can serve as a guideline to finding optimal operational parameters. By taking advantage of the flip protocol, the global lowest logical error and decoding time are obtained together for the presented parameters when – perhaps surprisingly – the lowest seepage rate is chosen. Combining numerical input with direct experimental optimization of complex tradeoff choices in hardware parameters, software aspects and decoding model can be expected to become increasingly more relevant, encompassing not only logical errors [domokos2024characterization, marton2024optimal] but other computational aspects such as decoding time, as above.
@@
Conclusion and outlook. The adaptive readout method the we propose can significantly reduce both logical errors and decoding time. The highest impact of the protocol is expected if the conditional or adapted gates on the check qubits could be implemented without delays. While current hardware supports conditional operation in real time, it requires a finite time to execute after inquiring the measurement outcome. As our protocol doesn’t require conditional operations immediately after measurement, in principle the whole duration of a syndrome cycle can be used to process and prepare the required instruction.
@@
Analyzing the expected tradeoffs between different mechanisms for suppression of errors in QEC settings could be a powerful tool to test global optimization approaches (and calibration approaches as in App. A). The simulation code used to produce the data for this work is available as open source [qec_flip_repo], and it can be extended and used to quantify more elaborate error models or conditional protocols. As another example for state-dependent noise, in App. B we probe preparation faults correlated to readout errors (i.e., a preparation error occurring conditioned on a readout error, as in the so-called “conditional reset” scheme that benefits from a fast implementation given the preceding measuring), seeing a dramatic increase in the logical error rates under certain conditions, with some reduction using the flip protocol.